OpenStudy (babynini):

Find all solutions for equation, find solutions in interval [0,2pi) sq3 tan 3(theta)+1=0

OpenStudy (babynini):

@peachpi the victory was short lived lol

OpenStudy (babynini):

3(theta)=arctan(-sq3/3) ?

OpenStudy (anonymous):

This is the original? \[\sqrt{3\tan 3\theta+1}=0\]

OpenStudy (babynini):

hm, no. Let me write it out for you :) \[\sqrt{3}\tan3(\theta)+1=0\]

OpenStudy (anonymous):

got you

OpenStudy (babynini):

\[\tan3(\theta)=\frac{ -\sqrt{3} }{ 3 }\]

OpenStudy (babynini):

right?

OpenStudy (anonymous):

yes

OpenStudy (babynini):

So then after quite a bit of simplifying theta = -0.26 ?

OpenStudy (babynini):

oh wait -0.17

OpenStudy (anonymous):

That one's on the unit circle, so they might be looking for an exact answer. -π/18, which is about -0.17, so you're right

OpenStudy (babynini):

pi/18 is on the unit circle? o.o

OpenStudy (anonymous):

Well pi/6 is.

OpenStudy (anonymous):

You had 3Θ = arctan (-1/√3) 3Θ = -π/6 Θ = -π/18

OpenStudy (babynini):

Why does 3(theta) = -pi/6 ?

OpenStudy (babynini):

sorry, these are probably really elementary questions but I want to make sure and understand.

OpenStudy (anonymous):

because arctan has a domain from -pi/2 to pi/2 (because tan positive in the 1st quadrant, negative in the 4th). The angle in that domain with a tan of -1/√3 is -π/6

OpenStudy (babynini):

hm..k.

OpenStudy (anonymous):

does that make sense?

OpenStudy (babynini):

Yeah I just hadn't made the connection so now i've got to think it through a bit :P

OpenStudy (babynini):

So now to find it in the right quadrants we add pi?

OpenStudy (anonymous):

ok so this one's a little trickier since the tan is negative. The restriction on the equation is 0 to 2pi, so we actually need to add 2pi to -pi/6 to get the 1st solution for 3pi

OpenStudy (anonymous):

We'll have 2 solutions for 3pi between 0 and 2pi: 11π/6 in the 4th quadrant (-pi/6 + 2π) 5π/6 in the 2nd quadrant (-π/6 + π)